The average age of Donald, his wife and their two children is 23 years. His wife is just 4 years younger than Donald himself and his wife was 24 yrs old when his daughter was born. He was 32 yrs old when his son was born. The average of Donald and his daughter is?
Answer by Edwin McCravy(20060)

The average age of Donald, his wife and their two children is 23 years. His
wife is just 4 years younger than Donald himself and his wife was 24 yrs old
when his daughter was born. He was 32 yrs old when his son was born. The
average of Donald and his daughter is?
>>...The average age of Donald, his wife and their two children is 23
years...<< 

Let Donald's age = x

>>...His wife is just 4 years younger than Donald himself...<<

His wife's age = Donald's age - 4 = x-4

>>...his wife was 24 yrs old when his daughter was born...<<

So his daughter is 24 years younger than his wife, so we get the
daughter's age by subtracting 24 years from the wife's age of x-4

daughter's age = mother's age minus 24 = 
x-4 - 24 which simplifies to x-28

>>...He was 32 yrs old when his son was born...<<

So his son is 32 years younger than Donald, so we get the son's 
age by subtracting 32 years from Donald's age of x.

son's age = Donald's age minus 32 = x-32

>>...The average age of Donald, his wife and their two children is 23 years...<<

If add up all their ages and divide by 4 we get 23, so

(Donald's age + Wife's age + daughters age + son's age) divided by
4 equals 23:

x + x-4 + x-28 + x-32
--------------------- = 23        
         4

              4x - 64
              ------- = 23
                 4

Multiply both sides by 4

              4x - 64 = 92

add 64 to both sides 

                   4x = 156

divide both sides by 4

                    x = 39

So Donald's age is 39

Donald's wife's age is x-4 or 39-4 or 35.

Their daughter's age is x-28 or 39-28 or 11

Their son's age is x-32 or 39-32 or 7


Therefore the average of Donald's and his daughter's ages is
found by adding their ages and dividing by 2:

39 + 11    50
------- = ---- = 25
  2        2

Edwin